# VGPT-PINN: Viscosity-enhanced Generative Pre-Trained Physics Informed Neural Networks for parameterized nonlinear conservation laws

### v1

@@ 5 Conclusion
In this paper, we develop the VGPT-PINN framework for the parameterized nonlinear conservation law problems and apply it to the inviscid Burgers’ equation and the Euler equations. Our approach extends the capabilities of traditional PINNs and the recently introduced GPT-PINN by incorporating nonlinear model reduction techniques and viscosity awareness while maintaining an unsupervised learning structure.
For Burgers’ equation, the VGPT-PINN uses only one or two neurons to handle the discontinuities inherent in the shock waves, demonstrating superior accuracy and efficiency compared to traditional methods. In the case of the Euler equations, the VGPT-PINN’s two-stage training process, with differentiated learning rates for the output and transform layers, ensured robust and rapid approximation of the solutions. This approach successfully managed the challenges posed by the shock waves and shock contact, with only 4 or 5 neurons, accurately capturing the dynamics of shock wave propagation and interaction. In both cases, the framework’s ability to adaptively learn the system’s parametric dependencies and incrementally expand the hidden layers proved crucial in capturing the complex behavior of shock interactions.

### v2

@@ 5 Conclusion
In this paper, we develop the EGPT-PINN framework for the parameterized nonlinear conservation law problems and apply it to the inviscid Burgers’ equation and the Euler equations. Our approach extends the capabilities of traditional PINNs and the recently introduced GPT-PINN by incorporating nonlinear model reduction techniques and entropy awareness while maintaining an unsupervised learning structure.
For Burgers’ equation, the EGPT-PINN uses only one or two neurons to handle the discontinuities inherent in the shock waves, demonstrating superior accuracy and efficiency compared to traditional methods. In the case of the Euler equations, the EGPT-PINN’s two-stage training process, with differentiated learning rates for the output and transform layers, ensured robust and rapid approximation of the solutions. This approach successfully managed the challenges posed by the shock waves and shock contact, with only 3 to 5 neurons, accurately capturing the dynamics of shock wave propagation and interaction. In both cases, the framework’s ability to adaptively learn the system’s parametric dependencies and incrementally expand the hidden layers proved crucial in capturing the complex behavior of shock interactions.